# Representation Shattering in Transformers: A Synthetic Study with Knowledge Editing

### v3

@@ Appendix A Setup Details
We will publicly release the source code for our work on GitHub at a later time.
A.1 Pseudo-Code
@@
1 function generateSequence()
2       $x_{p}\sim U(X)$ from a uniform distribution over the entities.
3       $S$ = [$x_{p}$]
4       entity_flag $\leftarrow$ False
       // Create a sequence of alternating entities and relations
5       while len($S$) < context_size do
6             if (entity_flag) then
                   // Add an entity that completes a valid fact
7                   Set $x_{n}$ such that $(x_{p},\vec{r},x_{n})$ is a fact in the knowledge graph $G$.
8                   $S$.append($x_{n}$)
9                   $x_{p}\leftarrow x_{n}$
10            else
                   // Add a composition of relations
11                   $K\sim U(\{1,2\})$. $\vec{r}$ = [ ]
12                   for ($i$ in 1 to $K$) do
13                         $r\sim U(R)$
14                         S.append($r$)
15                         $\vec{r}$.append($r$)
16
17                  Set $x_{n}$ such that $(x_{s},\vec{r},x_{n})$ is a fact in the knowledge graph $G$.
18
@@ Appendix B Data Generation Process Details
Additionally, only when generating the training dataset, we drop sequences which contain one direction of a pair of conjugate facts with fixed probability $p$. In other words, if the fact $(x_{i},r,x_{j})$ always implies that $(x_{j},r^{\prime},x_{i})$ is a valid fact (i.e. $r=$ I_C1 and $r^{\prime}=$ I_A1), one of $(x_{i},r,x_{j})$ or $(x_{j},r^{\prime},x_{i})$ may be restricted to inclusion in the training dataset by composition only (with probability $p$). Holding out these relations allows us to benchmark the model’s logical inference capabilities on relations it could not have directly memorized from the training dataset. In practice, we set the probability $p=\frac{2}{3}$.
Appendix C Model Architecture
@@
$\mathcal{L}(\mathbf{x}_{n},\text{target }n)=-\log\Big{(}\dfrac{\exp(\beta x_{n,\text{ target }n})}{\sum_{v=0}^{\#\text{tokens}}\exp(\beta x_{n,v})}\Big{)}=-\log\Big{(}\underbrace{\text{softmax}(\beta\mathbf{x}_{n})_{\text{target }n}}_{\text{prob}(\text{target }n)}\Big{)}$
Appendix D Knowledge Editing Methods
D.1 Rank-One Model Editing (ROME)
D.1.1 Algorithm Definition
Rank-One Model Editing (ROME), proposed by Meng et al. (2022a), is a popular knowledge editing algorithm used on LLMs. Their contributions are two-fold: first, through “causal tracing,” they find that early-layer MLP modules of transformer models are implicated in encoding factual associations. Second, interpreting feed-forward layers as linear associative memories encoding key-value pairs, ROME applies a rank-one update to the MLP weights.
@@ D.1.2 Implementation
D.2 Mass-Editing Memory in a Transformer (MEMIT)
D.2.1 Algorithm Definition
The Mass-Editing Memory in a Transformer (MEMIT) algorithm proposed by Meng et al. (2023) generalizes ROME (Meng et al., 2022a) to inject many factual associations at once. Rather than modifying a single layer for a single key-value pair, MEMIT identifies multiple MLP layers (via causal tracing) which encode facts. For each fact $(x_{i},r,x_{j}^{*})$, a small residual $\bm{\delta}_{i}$ is computed to increase $P(x_{j}^{*}\mid x_{i},r)$. These $\bm{\delta}_{i}$ vectors are then batched and distributed across the identified layers. In each layer $\ell$, let $\mathbf{K}$ and $\mathbf{M}$ respectively stack the new keys and target values derived from $\{\bm{\delta}_{i}\}$. Extending ROME, the weight update for layer $\ell$ is
@@ D.2.1 Algorithm Definition
D.2.2 Implementation
For our experiments with Llama 3 8B Instruct and Mamba 2.8B, we simply apply public implementations with off-the-shelf hyperparameters (Gupta et al., 2024a; b; Sharma et al., 2024).
Appendix E Visualization Methods
@@
Appendix F Additional Synthetic DGP Experiments
F.1 Alternative Editing Methods and Models
F.1.1 Model Accuracy
@@
F.1.2 Representation Shattering Metric
In Tab. 2, we showed that increasing the distance of the counterfactual edit results in an increase in the extent of shattering, as numerically captured by $R(D_{*})$. In similar spirit to Appx. F.1.1, we seek to verify whether this relationship between counterfactual edit distance and representation shattering holds for methods other than ROME, i.e. MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and PMET (Li et al., 2024). We present our results in Tab. 4.
@@ F.1.2 Representation Shattering Metric
Table 4: Results from Tab. 2, replicated using the alternative knowledge editing methods of MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and PMET (Li et al., 2024). These successors to ROME achieve lower amounts of representation shattering overall, coinciding with their more favorable performance in Appx. F.1.1. However, the relationship between greater counterfactual edit distance $d$ and greater representation shattering $R(D_{*})$ still robustly holds for all methods. This result again shows that various KE methods struggle in similar ways: specifically, the greater the inconsistency between the model’s original knowledge and the edited fact, the greater the resulting distortion upon the model’s representations.
F.2 Independence of Cyclic Orders
@@
F.3 Manifolds for Various Representation Extraction Points
@@ F.4 Manifolds for All Relations
F.5 Counterfactual Editing
F.5.1 Distribution of Degredations for Counterfactual Edits
@@
F.5.2 Additional Visualizations
In Fig. 6, we showcase an example of the change in accuracies and representation manifolds when applying a counterfactual edits (specifically for fact 1154.I_C1). For a more representative view, we additionally provide more examples of counterfactual edits (with both raw and pre-processed versions side-by-side, as described in Appx. E).
@@ F.5.2 Additional Visualizations
Appendix G Additional Naturalistic LLM Experiments
G.1 Knowledge Editing on Mamba
@@
Figure 15: Fig. 7, replicated using Mamba 2.8B (Gu & Dao, 2023). Like in Llama 3 8B Instruct, the ring structure shatters for larger counterfactual edit distances (activations were collected at the layer 38 output; see Appx. G.2 for other extraction points). This result demonstrates that our findings are not limited to specific models and can be extended to various architectures.
G.2 LLM Representation Extraction Points
To measure and visualize representation shattering in Sec. 4.5 and Appx. G.1, we extracted intermediate representations from Llama 3 8B Instruct and Mamba 2.8B at the output of layers 19 and 38, respectively. We chose these layers because they were the earliest points in each model for which cyclic representations could be observed at the last token position (as shown in Fig. 16 and Fig. 17).
@@ G.2 LLM Representation Extraction Points
G.3 Knowledge Editing with Naturalistic Trees

### v5

@@ Appendix A Setup Details
A.1 Source Code
[…]
A.2 Pseudo-Code
@@ Appendix B Data Generation Process Details
[…]
Appendix C Prompt Formats
  • Compositional Inference: The model is given a chain of relations (e.g., $x_{i}\,r_{1}\,r_{2}$) and must produce $x_{k}$, with $(x_{i},r_{1},x_{j})$ and $(x_{j},r_{2},x_{k})$ seen separately in training but not composed. Input: $\text{ctx}\,x_{i}\,\vec{r}$ where $\vec{r}=r_{1}r_{2}$.
  • Logical Inference: The model is asked about a fact $(x_{j},r^{\prime},x_{i})$ where $(x_{i},r,x_{j})$ was seen but $r^{\prime}$ was withheld. If the model understands relation symmetry, it can infer the inverse.
Appendix D Model Architecture
@@
$\mathcal{L}(\mathbf{x}_{n},\text{target }n)=-\log\Big{(}\dfrac{\exp\small(\beta x_{n,\text{ target }n}\small)}{\sum_{v=0}^{\#\text{tokens}}\exp\small(\beta x_{n,v}\small)}\Big{)}=-\log\Big{(}\underbrace{\text{softmax}(\beta\mathbf{x}_{n})_{\text{target }n}}_{\text{prob}(\text{target }n)}\Big{)}$
Appendix E Knowledge Editing Methods
E.1 Rank-One Model Editing (ROME)
E.1.1 Algorithm Definition
Rank-One Model Editing (ROME), proposed by Meng et al. (2022a), is a popular knowledge editing algorithm used on LLMs. Their contributions are two-fold: first, through “causal tracing,” they find that early and mid-layer MLP modules of transformer models are implicated in encoding factual associations (see Appx. G.2.1). Second, interpreting feed-forward layers as linear associative memories encoding key-value pairs, ROME applies a rank-one update to the MLP weights.
@@ E.1.2 Implementation
E.2 Mass-Editing Memory in a Transformer (MEMIT)
E.2.1 Algorithm Definition
The Mass-Editing Memory in a Transformer (MEMIT) algorithm proposed by Meng et al. (2023) generalizes ROME (Meng et al., 2022a) to inject many factual associations at once. Rather than modifying a single layer for a single key-value pair, MEMIT identifies multiple MLP layers (via causal tracing) which encode facts. For each fact $(x_{i},r,x_{j}^{*})$, a small residual $\boldsymbol{\delta}_{i}$ is computed to increase $P(x_{j}^{*}\mid x_{i},r)$. These $\boldsymbol{\delta}_{i}$ vectors are then batched and distributed across the identified layers. In each layer $\ell$, let $\mathbf{K}$ and $\mathbf{M}$ respectively stack the new keys and target values derived from $\{\boldsymbol{\delta}_{i}\}$. Extending ROME, the weight update for layer $\ell$ is
@@ E.2.1 Algorithm Definition
E.2.2 Implementation
For our experiments with Llama 3 8B Instruct and Mamba 2.8B, we simply apply public implementations with off-the-shelf hyperparameters (Gupta et al., 2024a, b; Sharma et al., 2024).
Appendix F Visualization Methods
@@
Appendix G Additional Synthetic DGP Experiments
G.1 Alternative Editing Methods and Models
G.1.1 Model Accuracy
@@
G.1.2 Representation Shattering Metric
In Tab. 2, we showed that increasing the distance of the counterfactual edit results in an increase in the extent of shattering, as numerically captured by $R(D_{*})$. In similar spirit to Appx. G.1.1, we seek to verify whether this relationship between counterfactual edit distance and representation shattering holds for methods other than ROME, i.e. MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and PMET (Li et al., 2024). We present our results in Tab. 4.
@@ G.1.2 Representation Shattering Metric
Table 4: Results from Tab. 2, replicated using the alternative knowledge editing methods of MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and PMET (Li et al., 2024). These successors to ROME achieve lower amounts of representation shattering overall, coinciding with their more favorable performance in Appx. G.1.1. However, the relationship between greater counterfactual edit distance $d$ and greater representation shattering $R(D_{*})$ still robustly holds for all methods. This result again shows that various KE methods struggle in similar ways: specifically, the greater the inconsistency between the model’s original knowledge and the edited fact, the greater the resulting distortion upon the model’s representations.
G.2 Validation of KE Algorithms’ Assumptions
The knowledge editing methods we test across our experiments—namely ROME (Meng et al., 2022a), MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and PMET (Li et al., 2024)—make assumptions about the mechanisms by which language models store factual associations in their learned weights. Here, we verify those assumptions to ensure that our application of KE to our synthetic experimental setup is in fact appropriate.
G.2.1 Causal Tracing of Factual Associations
ROME (Meng et al., 2022a) and its follow-up derivatives rely on “causal tracing” to locate facts within the parameters of large pretrained autoregressive transformer models. This is essentially causal mediation analysis applied to the internal states of transformer models, with scores attributing each state’s contribution toward a correct factual prediction. KE methods build upon the finding that, early/middle MLP layers play a decisive role in recalling key-value facts about the subject (Meng et al., 2022a). Below, we investigate whether this finding also generalizes to our toy model by performing causal tracing.
Figure 9: The average “Indirect Effect” of MLP output states (for the last subject token) on the final fact output elicited from the model, computed across $100$ randomly sampled prompts. While our model only has two layers, causal tracing attributes higher scores to the first MLP layer rather than the second. This aligns with the findings of Meng et al. (2022a) who showed that early and mid-layer MLPs play a larger role in factual recall than later sites. We therefore target the first layer MLPs for our knowledge editing experiments.
G.2.2 Transformer MLPs as Linear Associative Memory
Prior works view the $\mathbf{W}_{proj}^{(l)}$ matrix of the MLP layer within the Transformer architecture as linear associative memory. Under this perspective, the MLP layers act as key-value stores for vector keys and corresponding vector values (Kohonen, 2009; Anderson, 1972; Geva et al., 2020; Bau et al., 2020). Below, we qualitatively investigate whether this observation also holds for facts learned by our toy model.
Refer to caption Refer to caption Refer to caption
(a.i) Key (b.i) Key (c.i) Key
Refer to caption Refer to caption Refer to caption
(a.ii) Value (b.ii) Value (c.ii) Value
(a) 420.I_C2=15 (b) 1650.I_C3=15 (c) 1234.II_C2=2007
Figure 10: Inputs (keys) and outputs (values) of the layer $1$ MLP module, projected to 2D with PCA. Gray corresponds to randomly sampled prompts across all possible facts, red to randomly sampled prompts for a specific fact, and blue to the average position of the red points. Comparing (a) and (b), we find that fact prompts which resolve to the same final target entity in the same subgraph (i.e., entity 15)
share similar keys and values. Meanwhile, a completely different fact (like 1234.II_C2=2007) is stored as a clearly distinct key-value pair. These patterns support the key-value framework as compatible with our synthetic setup.
G.3 Independence of Cyclic Orders
@@
G.4 Manifolds for Various Representation Extraction Points
@@ G.5 Manifolds for All Relations
G.6 Counterfactual Editing
G.6.1 Distribution of Degredations for Counterfactual Edits
@@
G.6.2 Additional Visualizations
In Fig. 6, we showcase an example of the change in accuracies and representation manifolds when applying a counterfactual edits (specifically for fact 1154.I_C1). For a more representative view, we additionally provide more examples of counterfactual edits (with both raw and pre-processed versions side-by-side, as described in Appx. F).
@@ G.6.2 Additional Visualizations
Appendix H Additional Naturalistic LLM Experiments
H.1 Knowledge Editing on Mamba
@@
Figure 17: Fig. 7, replicated using Mamba 2.8B (Gu & Dao, 2023). Like in Llama 3 8B Instruct, the ring structure shatters for larger counterfactual edit distances (activations were collected at the layer 38 output; see Appx. H.2 for other extraction points). This result demonstrates that our findings are not limited to specific models and can be extended to various architectures.
H.2 LLM Representation Extraction Points
To measure and visualize representation shattering in Sec. 4.5 and Appx. H.1, we extracted intermediate representations from Llama 3 8B Instruct and Mamba 2.8B at the output of layers 19 and 38, respectively. We chose these layers because they were the earliest points in each model for which cyclic representations could be observed at the last token position (as shown in Fig. 18 and Fig. 19).
@@ H.2 LLM Representation Extraction Points
H.3 Knowledge Editing with Naturalistic Trees